6.5.4.1 Consider the following Boolean function. It is required to implement it using 8x1 MUX and
4x1 MUX
F(x, y, z) = x' y' z' + x' y z' + x y' z + x y z
Answer #1

Solution:

F(x, y, z) = x' y' z' + x' y z' + x y' z + x y z

Now we have to implement the above Boolean function uisng 8×1 MUX and 4×1 MUX.
Answer #2

Given Boolean Function:

F(x,y,z)=xyz+xyz+xyz+xyz

Step 1: Simplify the Function (Optional)

First, check if the function can be simplified:

F(x,y,z)=xyz+xyz+xyz+xyz

Factor out common terms:

=z(xy+xy)+z(xy+xy)=z(x(y+y))+z(x(y+y))

Since y+y=1:

F(x,y,z)=zx+zx

This simplifies to:

F(x,y,z)=xz

 


Option 1: Implement Using an 8×1 MUX

An 8×1 MUX has 3 select lines (for 8 inputs) and can directly implement any 3-variable function.


  1. Truth Table for F(x,y,z):

    xyzF
    000 1(xyz)
    001 0
    010 1(xyz)
    011 0
    100 0
    101(xyz)
    1100
    1111(xyz)
  2. MUX Configuration:

    • Select Lines: S2=xS1=yS0=z.

    • Input Lines (D0–D7):

      D0=1,D1=0,D2=1,D3=0,D4=0,D5=1,D6=0,D7=1



Option 2: Implement Using a 4×1 MUX

4×1 MUX has 2 select lines, so we must reduce the function to 2 variables by fixing one variable as a select line.

  1. Choose x and y as Select Lines (S1=xS0=y):

    • The remaining variable z determines the MUX inputs.


  2. Express F in terms of z for each (x,y) combination:

    • x=0,y=0: F=z

    • x=0,y=1: F=z

    • x=1,y=0: F=z

    • x=1,y=1: F=z

  3. MUX Configuration:

    • D0 and D1: Connect to z (NOT z).

    • D2 and D3: Connect directly to z.

    • Input Lines (D0–D3):

      D0=z,D1=z,D2=z,D3=z

    • Logic for Inputs:


Answer:

  • 8×1 MUX: Directly implement using the truth table (inputs = 1, 0, 1, 0, 0, 1, 0, 1).

  • 4×1 MUX: Use x,y as select lines and z/z as inputs.


answered by: anonymous